Egbert invested $5,000 for one year, part at 9% annual interest and the rest at 12% annual interest. The interest at 9% earned $198 more than the interest from the investment at 12%. How much was invested at 9%?

Answer by mananth(16949)

Part I 9.00% per annum ------------- Amount invested =x
Part II 12.00% per annum ------------ Amount invested = y
5000
Interest----- 198.00

Part I 9.00% per annum ---x
Part II 12.00% per annum ---y
Total investment
x + y= 5000 -------------1
Interest on both investments
9.00% x + -12.00% y= 198
Multiply by 100
9 x + -12 y= 19800.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -45000.00
Add this to (2)
0 x -21 y= -25200
divide by -21
y = 1200
x= 5000-1200
=3800 at 9%